# Multiclass Portfolio Optimization via Variational Quantum Eigensolver with Dicke State Ansatz

### v2

@@ I Introduction
Quantum Computing (QC) represents a groundbreaking technology poised to solve problems that lie beyond the capabilities of classical computers. It operates by manipulating quantum systems and harnessing unique properties such as superposition, interference, and entanglement. This approach enables the execution of advanced algorithms designed to address complex challenges with remarkable efficiency. The potential of quantum computing spans multiple fields, promising significant impacts across industries and scientific disciplines; some of them include: quantum chemistry and material science [1, 2, 3], machine learning [4, 5, 6, 7, 8, 9], finance [10, 11, 12, 13, 14, 15, 16], and optimization [17, 18, 19, 20, 21, 22, 23].
@@
Optimization plays a fundamental role across various domains, with classic examples including the traveling salesman problem [24], vehicle routing [25], bin packing [26], and portfolio optimization [27]—the central focus of this article. These problems are traditionally tackled using classical techniques such as mixed-integer programming [28], approximation algorithms and metaheuristics [29], and neural networks [30, 31]. However, due to their computational complexity, these methods tend to have limitations in either their running time or solution quality as the number of variables increases.
@@
Portfolio optimization exemplifies this challenge, especially in multiclass scenarios that require allocation across diverse asset categories (e.g., equities, bonds, commodities). While classical approaches like the Markowitz mean-variance model [32] perform well for small-scale instances, multiclass formulations introduce combinatorial and equality constraints that lead to mixed-integer quadratic programs (MIQPs), which are hard to solve optimally, particularly in real-time or resource-limited settings.
@@
In response to these limitations, Variational Quantum Algorithms (VQAs) [33, 34] have emerged as promising tools for near-term quantum advantage. Designed for Noisy Intermediate-Scale Quantum (NISQ) devices, VQAs—especially the Variational Quantum Eigensolver (VQE) [35]—have demonstrated potential in solving combinatorial problems by mapping them to Ising Hamiltonians and approximating the ground state via parameterized quantum circuits. The effectiveness of this approach heavily depends on the choice of ansatz, particularly when handling complex constraints such as those in multiclass portfolio optimization.
To that end, this paper proposes the use of a Dicke state-based ansatz within the VQE framework to efficiently handle multiclass portfolio optimization, incorporating a realistic feature in this problem related to diversification of the investment. Our contributions are threefold: (i) we formulate a multiclass portfolio optimization problem suitable for quantum encoding using a parametrized Dicke state; (ii) we implement and benchmark the performance of the VQE using this ansatz on simulated quantum hardware; and (iii) evaluate the effect of classical optimizers in the convergence to the optimal solution.
@@
This paper is structured as follows: in Section II we introduce the problem of multiclass portifolio optimization, Section III introduces VQE and the standard ansatz. Section IV describes the problem formulation, the Dicke state ansatz, and the methodology adopted. Section V presents the numerical results, empirical findings, and discussions. Finally, Section VI summarizes the main contributions and outlines directions for future research in quantum finance.
@@ II Multiclass portfolio optimization
Figure 1: Illustrative example of a multiclass portfolio optimization problem with a predefined asset class allocation to ensure diversification aiming the reduction of market risk. The total number of assets considered is 820, distributed as follows: 500 stocks, 200 cryptocurrencies, 30 commodities, 10 ETFs, 60 REITs and 20 Bonds. For each class there are $C_{n,k}$ possible portfolios, where $n$ represents the total number of assets in the class and $k$ is the predefined number of assets to be selected. The goal of this portfolio optimization is to find the best set of assets that will produce a portfolio that satisfies the constraints maximizing the return and minimizing the risk.
Diversification reduces exposure to specific market segments, enhances portfolio resilience, and aligns with best practices in risk management. In its classical form, the objective is to select a subset of assets that minimizes portfolio risk while maximizing expected return, under a given level of risk aversion [32], as formulated below
@@
where $x$ is a binary decision vector such that an entry equal to 1 indicates the inclusion of a corresponding asset in the portfolio, $\Sigma$ is the covariance matrix of asset returns, capturing the portfolio’s overall risk, while $\mu$ represents the expected return vector. The parameter $r_{f}$ denotes the risk-free rate, such as the return of US Treasury bonds. The matrix $A_{m\times n}$ encodes the linear constraints that govern portfolio selection, where $m$ is the number of constraint equations and typically corresponds to the number of asset classes or diversification. To simplify and reduce the number of experiment parameters, we set $q=0.5$.
@@ III Variational Quantum Algorithms and Ansatz
The primary VQAs for combinatorial optimization are VQE [35] and QAOA [36]. In this work, we focus only on VQE in the context of multiclass portfolio optimization (see Fig.2). Both approaches utilize a parameterized quantum circuit, commonly referred to as an ansatz, together with a classical optimization routine. Their objective is to minimize the expectation value of the problem Hamiltonian ($H$) which encodes the specific combinatorial optimization problem to be solved, as showed in the equation below:
@@
The expectation value defined in equation (2) is lower bounded by the minimum eigenvalue of the Hamiltonian $H$, known as the energy of the ground state $E_{0}$. The primary goal of this approach is to identify an appropriate ansatz along with an optimal set of parameters $\vec{\theta^{*}}$, such that the expected value computed $\langle\psi(\vec{\theta^{*}})|H|\psi(\vec{\theta^{*}})\rangle$ is equal or close to $E_{0}$. Quantum mechanics guarantees the existence of this lower bound [37], although its exact value is generally unknown beforehand. Thus, the variational method provides a practical way to approximate both the ground-state wavefunction and its corresponding energy.
Figure 2: An illustrative example of VQE routine, which is defined by a quantum state preparation conducted on a quantum device or simulator, followed by an optimization process in a classical computer. In the quantum routine, we start with a quantum state where all qubits are in state $|0\rangle$, this initial state evolves according to the unitary $U(\vec{\theta_{i}})$, which defines ansatz structure and receives a set of parameters that will define the states probability distribution extracted from a certain number of measurements. The classical routine is focused on updating the set of parameters $\vec{\theta_{i}}$, in order to minimize the expectation value of the Hamiltonian that encodes the optimization problem. This process is repeated until the maximum number of iterations or when other stopping criteria is achieved.
The main difference between VQE and QAOA is that the latter is a special case of the former, once the QAOA ansatz has a well-defined structure based on the adiabatic theorem [37], which, in general, leads to deeper quantum circuits. There is a trade-off between these algorithms, VQE offers a shallow quantum circuit with a higher number of parameters, meanwhile QAOA offers a deeper ansatz with $2p$ parameters, where $p$ represents the number of layers, so in some cases QAOA could solve a problem using fewer parameters but at the cost of using a deeper circuit. In any case, both have the potential to extract useful results from NISQ devices.
Some key challenges in applying VQA to practical scenarios is identifying the optimal ansatz, determining efficient parameter initialization methods, and selecting the most suitable classical optimizer. In this study, we systematically explored various ansatzes to determine the most suitable configuration for our specific problem. The investigated ansatzes can be categorized into three distinct types: $(i)$ simple $R_{y}$ rotation gates, $(ii)$ the extensively studied TwoLocal ansatz along with its variants [13], and $(iii)$ the parameterized Dicke state. A detailed schematic representation of each ansatz is provided in Fig.3.
@@ IV Portfolio Optimization and Dicke State
To address portfolio optimization with quantum computing, first we need to convert Eq.(1) to a Quadratic Unconstrained Binary Optimization (QUBO) problem [38], then the QUBO is converted to an Ising model through the change of variables $x_{i}=\frac{(1-z_{i})}{2}$ [39]
@@
[41] $O(k\log\frac{n}{k})$ $O(nk)$ N/A all-to-all
[41] $O(k\sqrt{\frac{n}{k}})$ $O(nk)$ N/A grid
[42] 2(nk)2(n-k) $2nk-3k^{2}$ $nk-\frac{3k^{2}}{2}$ LNN
[42] $2n$ $nk-\frac{k^{2}}{2}$ $\frac{n(k+1)}{2}-\frac{k^{2}}{4}$ LNN
@@
To enhance the approach, we can leverage knowledge about the optimization problem structure to design more efficient circuits. For example, let us consider a portfolio optimization problem with $n$ variables and constraints that specify the exact number of $k$ assets that must be selected to minimize risk and maximize return. In this situation, feasible candidate solutions must have a Hamming weight equal to $k$. The set of feasible solutions is given by $C_{n,k}=n!/k!(n-k)!$. Consequently, among all $2^{n}$ quantum states, we eliminate those that do not meet the specified constraints. By initializing the quantum state in a superposition that encompasses only feasible solutions, the search space is effectively minimized from $O(2^{n})$ to $O(n^{k})$. Hence, the quantum state suitable for handling such constraints is known as the Dicke state [42], which is a quantum state related to a fundamental model of quantum optics that describes the interaction between light and matter [43]. The general formula for a uniform distribution of quantum states with $n$ qubits and Hamming weight $k$ is defined by [44]
$|D_{k}^{n}\rangle=\binom{n}{k}^{-1/2}\sum_{i}\mathcal{P}_{i}|0\rangle^{\otimes\ (n-k)}\otimes|1\rangle^{\otimes\ k},$ (4)
where $\mathcal{P}_{i}$ represents each possible permutation of a quantum state with $n$ qubits with $k$ qubits equal $|1\rangle$. There are several different implementations of the Dicke state (see Table 1). The quantum state defined by Eq.(4), was used as the initial state for QAOA in [45, 46, 47, 48, 49], and in these references the authors tested a variety of mixers. Beyond the scope of optimization, Dicke states are relevant to the following fields: quantum game theory [50], quantum networks [51], quantum metrology [52], quantum error correction [53] and quantum storage [54].
@@
where $a_{i}(\vec{\theta})$ represents amplitude probability as a function of the parameters $\vec{\theta}$. A similar approach was presented by [42], which focused solely on the preparation of a single Dicke state, thus mimicking portfolio optimization without diversification constraints. In this paper, we use the implementation given by [40], creating a non-uniform Dicke state parameterizing the circuit implementation and incorporating the multiclass optimization by including multiple Dicke states representing the multiple classes. Each Dicke state represents a class with a set of products (number of qubits $n$) and will encode the constraints of the exact number of assets which will be selected by class (parameter $k$) (see Fig.3a). The equation below dictates the number of parameters of a Dicke state ansatz and it was derived empirically, for further details see Appendix XII
$n_{p}=\sum_{i=1}^{m}k_{i}n_{i}-\frac{k_{i}(k_{i}+1)}{2}.$ (6)
where $k_{i}$ is the number of states, $|1\rangle$, that corresponds to the Hamming weight associated with the budget constraint. For $m=1$, we have a unique Dicke state ansatz with $n$ qubits and Hamming weight $k$, but $m>1$ implies a tensor product of different Dicke states. The summation is over the number of classes represented by the number of Dicke states in the tensor product of the initialization.
@@ V Results
We addressed the portfolio optimization problem using the SCIP optimizer, treating its solution as the benchmark reference (see AppendixX, Table 4). The running time was determined by averaging the results from $100$ executions. This average running time served as the basis for comparing the performance of classical methods with hybrid VQE routines. However, it is important to note that we did not expect that the VQE approach would outperform the classical methods in terms of speed.
The experiments were performed in three different scenarios as described in Table 2. For Scenario I, we ran the VQE algorithm for $20$ different ansatzes (Dicke state, $R_{y}$, and $18$ variations of Two Local). For Scenarios II and III only the Dicke state was used. For all scenarios, we tested $5$ classical optimizers with $1000$ iterations, $100$ randomly sampled initial points (ansatz parameters), $4096$ shots per circuit, totalizing more than $60$ billion executions ($\approx n_{ansatz}\times n_{optimizers}\times n_{executions}\times n_{shots}\times n_{iterations}$) . More details of the variations of the two-local ansatz used here can be found in Appendix XI, Table 5. All data used in this work are publicly available and were obtained using the Yahoo Finance API.
@@
Scenario $n_{a}$ $n_{c}$ $n_{s}$ Ansatz $n_{p}$
I $10$ 11 44 $|D_{4}^{10}\rangle$ $30$
II $25$ 55 55 $|D_{1}^{5}\rangle^{\otimes\ 5}$ 2020
III $25$ 55 99 $|D_{2}^{5}\rangle^{\otimes\ 2}|D_{1}^{5}\rangle^{\otimes 2}|D_{3}^{5}\rangle$ $31$
Table 2: The table shows the different scenarios we use to evaluate the performance of parametrized Dicke state in the context of multiclass portifolio optimization. Where $n_{a}$ represents the number of assets, $n_{c}$ number of classes, $n_{s}$ amount of select assets, $n_{p}$ number of parameters of Dicke state ansatz. $|D_{k}^{n}\rangle$ is a superposition of all states with $n$ qubits and $k$ qubits equal $|1\rangle$. Each state represents a class where $n$ is the total number of assets available to choose and $k$ is the number of assets we must select from it. The parameter $\vec{\theta}$ was omitted in the state for better visualization.
@@
Figure 5: Plots $(a)$, $(b)$, and $(c)$ compare the approximation ratio ($a_{r}$) distributions before (gray) and after (colored) optimization for each scenario (columns) and optimizers (rows). The gray represent $a_{r}$ distributions with the initial parameters, while the colored ones reflect optimized parameters, based on $100$ experiments per optimizer. A vertical black dashed line marks $a_{r}=0.9$. Across all optimizers and scenarios, the initial distributions shift toward $a_{r}\geq 0.9$, indicating the positive impact of the hybrid approach and optimizers.In Scenario I (blue) all the optimizers guide the solution to the optimal region. In
@@
Another challenge in the evaluation of hybrid algorithms is to measure the effect of the classical optimizer on the solution. As mentioned before, we tested five different optimizers: CMA-ES [55], COBYLA [56], Random Sampler, SPSA [57] and QNSPSA [58]. The evaluation takes into account three different metrics: the approximation ratio, the frequency with which the right answer appears in the $100$ trials per optimizer, and the quality of the quantum output for each trial (the probability of measuring the target state). The approximation ratio is defined as [59, 60, 61, 22]
@@
where $\langle H\rangle_{\psi(\vec{\theta^{*})}}$ is the expected value of the problem Hamiltonian computed with the ansatz and the set of parameters obtained at the end of optimization. $E_{0}$ is the ground-state energy associated with the optimal solution that was calculated by computing the lowest eigenvalue for the problem Hamiltonian. The metric above measures how close or distant the solution is from the optimal. For example, an approximation ratio equal to $1$ means that the solution is equal to the optimal one.

### v3

@@ I Introduction
Quantum Computing (QC) represents a groundbreaking technology poised to solve problems that lie beyond the capabilities of classical computers. It operates by manipulating quantum systems and harnessing unique properties such as superposition, interference, and entanglement. This approach enables the execution of advanced algorithms designed to address complex challenges with remarkable efficiency. The potential of quantum computing spans multiple fields, promising significant impacts across industries and scientific disciplines; some of them include: quantum chemistry and material science [mcardle2020quantum, bauer2020quantum, von2021quantum], machine learning [biamonte2017quantum, liu2018quantum, schuld2019quantum, cong2019quantum, cerezo2022challenges, jerbi2023quantum], finance [egger2020quantum, ramos2021quantum, mugel2022dynamic, buonaiuto2023best, herman2023quantum, wilkens2023quantum, naik2025portfolio, Thakkar2024], and optimization [jarret2018improved, campbell2019applying, montanaro2020quantum, egger2021warm, magann2022feedback, abbas2024challenges, finvzgar2024quantum].
@@
Optimization plays a fundamental role across various domains, with classic examples including the traveling salesman problem [junger1995traveling], vehicle routing [toth2002vehicle], bin packing [martello2000three], and portfolio optimization [mansini2015linear]—the central focus of this article. These problems are traditionally tackled using classical techniques such as mixed-integer programming [junger200950], approximation algorithms and metaheuristics [gonzalez2007handbook], and neural networks [smith1999neural, prates2019learning]. However, due to their computational complexity, these methods tend to have limitations in either their running time or solution quality as the number of variables increases.
@@
Portfolio optimization exemplifies this challenge, especially in multiclass scenarios that require allocation across diverse asset categories (e.g., equities, bonds, commodities). While classical approaches like the Markowitz mean-variance model [markowitz1952portfolio] perform well for small-scale instances, multiclass formulations introduce combinatorial and equality constraints that lead to mixed-integer quadratic programs (MIQPs), which are hard to solve optimally, particularly in real-time or resource-limited settings.
@@
In response to these limitations, Variational Quantum Algorithms (VQAs) [cerezo2021variational, preskill2018quantum] have emerged as promising tools for near-term quantum advantage. Designed for Noisy Intermediate-Scale Quantum (NISQ) devices, VQAs, especially the Variational Quantum Eigensolver (VQE) [vqe2014], have demonstrated potential in solving combinatorial problems by mapping them to Ising Hamiltonians and approximating the ground state via parameterized quantum circuits. The effectiveness of this approach heavily depends on the choice of ansatz, particularly when handling complex constraints such as those in multiclass portfolio optimization.
To that end, this paper proposes the use of a Dicke state-based ansatz within the VQE framework to efficiently handle multiclass portfolio optimization, incorporating a realistic feature in this problem related to diversification of the investment. Our contributions are threefold: (i) Modeling advantage — we formulate a multiclass portfolio optimization problem suitable for quantum encoding by introducing a parameterized Dicke-state ansatz, in which diversification constraints are inherently satisfied through state preparation, eliminating the need for penalty calibration; (ii) Search-space advantage — by employing the Dicke state within the VQE framework, we drastically reduce the effective search space, restricting the sampling and optimization to the feasible manifold, which improves sample efficiency and convergence stability; and (iii) Empirical advantage — in our simulations, the Dicke ansatz combined with CMA-ES achieved higher approximation ratios and more frequent identification of the global optimum than standard ansatzes at comparable parameter counts. We make no claim of asymptotic quantum speedup but demonstrate meaningful structural and practical improvements within the variational paradigm.
@@
This paper is organized as follows: the first section introduces the problem of multiclass portfolio optimization. Subsequently, we provide an overview of VQE and the standard ansatz. We then describe the problem formulation, the Dicke state ansatz, and the methodology employed. Finally, we present the numerical results, empirical findings, and discussions. The paper concludes with a summary of the main contributions and outlines directions for future research in quantum finance.
@@ II Multiclass portfolio optimization
Figure 1: Illustrative example of a multiclass portfolio optimization problem with a predefined asset class allocation to ensure diversification aiming the reduction of market risk. The total number of assets considered is $820$, distributed as follows: $500$ stocks, $200$ cryptocurrencies, $30$ commodities, $10$ ETFs, $60$ REITs and 2020 Bonds. For each class there are $C_{n,k}$ possible portfolios, where $n$ represents the total number of assets in the class and $k$ is the predefined number of assets to be selected. The goal of this portfolio optimization is to find the best set of assets that will produce a portfolio that satisfies the constraints maximizing the return and minimizing the risk.
Diversification reduces exposure to specific market segments, enhances portfolio resilience, and aligns with best practices in risk management. In its classical form, the objective is to select a subset of assets that minimizes portfolio risk while maximizing expected return, under a given level of risk aversion [markowitz1952portfolio], as formulated below
@@
where $x$ is a binary decision vector such that an entry equal to 1 indicates the inclusion of a corresponding asset in the portfolio, $\Sigma$ is the covariance matrix of asset returns, capturing the portfolio’s overall risk, while $\mu$ represents the expected return vector. The parameter $r_{f}$ denotes the risk-free rate, such as the return of US Treasury bonds. The matrix $A_{m\times n}$ encodes the linear constraints that govern portfolio selection, where $m$ is the number of constraint equations and typically corresponds to the number of asset classes or diversification. In the example of Fig. 1, $a_{ij}=1$ whenever asset $j$ belongs to class $i$ and $a_{ij}=0$ otherwise. To simplify and reduce the number of experiment parameters, we set $q=0.5$.
@@ III Variational Quantum Algorithms and Ansatz
The primary VQAs for combinatorial optimization are VQE [vqe2014] and the Quantum Approximate Optimization Algorithm (QAOA) [farhi2014quantum]. In this work, we focus only on VQE in the context of multiclass portfolio optimization (see Fig.2). Both approaches utilize a parameterized quantum circuit, commonly referred to as an ansatz, together with a classical optimization routine. Their objective is to minimize the expectation value of the problem Hamiltonian ($H$) which encodes the specific combinatorial optimization problem to be solved, as showed in the equation below:
@@
The expectation value defined in equation (2) is lower bounded by the minimum eigenvalue of the Hamiltonian $H$, known as the energy of the ground state $E_{0}$. The primary goal of this approach is to identify an appropriate ansatz along with an optimal set of parameters $\vec{\theta^{*}}$, such that the expected value computed $\langle\psi(\vec{\theta^{*}})|H|\psi(\vec{\theta^{*}})\rangle$ is equal or close to $E_{0}$. Quantum mechanics guarantees the existence of this lower bound [griffiths2019introduction], although its exact value is generally unknown beforehand. Thus, the variational method provides a practical way to approximate both the ground-state wavefunction and its corresponding energy.
Figure 2: An illustrative example of a VQE routine, which is defined by a quantum state preparation conducted on a quantum device or simulator, followed by an optimization process in a classical computer. In the quantum routine, we start with a quantum state where all qubits are in state $|0\rangle$, this initial state evolves according to the unitary $U(\vec{\theta_{i}})$, which defines ansatz structure and receives a set of parameters that will define the states probability distribution extracted from a certain number of measurements. The classical routine is focused on updating the set of parameters $\vec{\theta_{i}}$, in order to minimize the expectation value of the Hamiltonian that encodes the optimization problem. This process is repeated until the maximum number of iterations or when other stopping criteria are achieved.
The main difference between VQE and QAOA is that the latter is a special case of the former, once the QAOA ansatz has a well-defined structure based on the adiabatic theorem [griffiths2019introduction], which, in general, leads to deeper quantum circuits. There is a trade-off between these algorithms, VQE offers a shallow quantum circuit with a higher number of parameters, meanwhile QAOA offers a deeper ansatz with $2p$ parameters, where $p$ represents the number of layers, so in some cases QAOA could solve a problem using fewer parameters but at the cost of using a deeper circuit. In any case, both have the potential to extract useful results from NISQ devices.
Some key challenges in applying VQE to practical scenarios are identifying the optimal ansatz, determining efficient parameter initialization methods, and selecting the most suitable classical optimizer. In this study, we systematically explored various ansatzes to determine the most suitable configuration for our specific problem. The investigated ansatzes can be categorized into three distinct types: $(i)$ simple $R_{y}$ rotation gates, $(ii)$ the extensively studied TwoLocal ansatz along with its variants [buonaiuto2023best], and $(iii)$ the parameterized Dicke state. A detailed schematic representation of each ansatz is provided in Fig.3.
@@ IV Portfolio Optimization and Dicke State
To address portfolio optimization with quantum computing, first we need to convert Eq.(1) to a Quadratic Unconstrained Binary Optimization (QUBO) problem [glover2019quantum], then the QUBO is converted to an Ising model through the change of variables $x_{i}=\frac{(1-z_{i})}{2}$ [lucas2014ising]
@@
[bartschi2022short] $O(k\log\frac{n}{k})$ $O(nk)$ N/A all-to-all
[bartschi2022short] $O(k\sqrt{\frac{n}{k}})$ $O(nk)$ N/A grid
[wang2024variational] 2(nk)2(n-k) $2nk-3k^{2}$ $nk-\frac{3k^{2}}{2}$ LNN
[wang2024variational] $2n$ $nk-\frac{k^{2}}{2}$ $\frac{n(k+1)}{2}-\frac{k^{2}}{4}$ LNN
@@
To enhance the approach, we can leverage knowledge about the optimization problem structure to design more efficient circuits. For example, let us consider a portfolio optimization problem with $n$ variables and constraints that specify the exact number of $k$ assets that must be selected to minimize risk and maximize return. In realistic scenarios, usually $k\ll n$ in order to obtain simpler and more explainable portfolios with an adequate level of diversification without reducing potential returns. In this situation, feasible candidate solutions must have a Hamming weight equal to $k$. The size of the set of feasible solutions is given by $C_{n,k}=n!/k!(n-k)!$. Consequently, among all $2^{n}$ quantum states, we eliminate those that do not meet the specified constraints. By initializing the quantum state in a superposition that encompasses only feasible solutions, the search space is effectively reduced from $O(2^{n})$ to $O(n^{k})$. Hence, the quantum state suitable for handling such constraints is known as the Dicke state [wang2024variational], which is a quantum state related to a fundamental model of quantum optics that describes the interaction between light and matter [garraway2011dicke]. The general formula for a uniform distribution of quantum states with $n$ qubits and Hamming weight $k$ is defined by [bartschi2019deterministic]
$|D_{k}^{n}\rangle=C_{n,k}^{-1/2}\sum_{i}\mathcal{P}_{i}|0\rangle^{\otimes\ (n-k)}\otimes|1\rangle^{\otimes\ k},$ (4)
where $\mathcal{P}_{i}$ represents each possible permutation of a quantum state with $n$ qubits with $k$ qubits equal to $|1\rangle$. There are several different implementations of the Dicke state (see Table 1). The quantum state defined by Eq.(4), was used as the initial state for QAOA in [cook2020quantum, bartschi2020grover, brandhofer2022benchmarking, he2023alignment, niroula2022constrained], and in these references the authors tested a variety of mixers. Beyond the scope of optimization, Dicke states are relevant to the following fields: quantum game theory [ozdemir2007necessary], quantum networks [prevedel2009experimental], quantum metrology [toth2012multipartite], quantum error correction [ouyang2021permutation] and quantum storage [ouyang2021quantum].
Beyond their broad applicability, Dicke states also serve as a foundation for our proposed ansatz in variational quantum algorithms. In this study, we assume all-to-all qubit connectivity to emphasize the conceptual contribution of the Dicke-state formulation and its ability to enforce diversification constraints without penalty terms. Nevertheless, the approach is not limited to this topology. As shown in Table 1, existing Dicke-state preparation circuits support all-to-all, grid, and linear-nearest-neighbor (LNN) architectures, with different trade-offs in depth and two-qubit gate count. Hence, the framework is hardware-agnostic, as preparing each class subspace only requires initializing a fixed-Hamming-weight superposition.
@@
where $a_{i}(\vec{\theta})$ represents amplitude probability as a function of the parameters $\vec{\theta}$. A similar approach was presented by [wang2024variational], which focused solely on the preparation of a single Dicke state, thus mimicking portfolio optimization without diversification constraints. In this paper, we use the implementation given by [mukherjee2020preparing], creating a non-uniform Dicke state parameterizing the circuit implementation and incorporating the multiclass optimization by including multiple Dicke states representing the multiple classes. Each Dicke state represents a class with a set of products (number of qubits $n$) and will encode the constraints of the exact number of assets which will be selected by class (parameter $k$) (see Fig.3a). The equation below dictates the number of parameters $n_{p}$ of a Dicke state ansatz
$n_{p}=\sum_{i=1}^{m}k_{i}n_{i}-\frac{k_{i}(k_{i}+1)}{2}$ (6)
where $k_{i}$ is the number of states $|1\rangle$, that corresponds to the Hamming weight associated with the budget constraint. Equation (6) was derived empirically, for further details see Suplementary Material.
For $m=1$, we have a unique Dicke state ansatz with $n$ qubits and Hamming weight $k$, but $m>1$ implies a tensor product of different Dicke states. The summation is over the number of classes represented by the number of Dicke states in the tensor product of the initialization.
@@ V Results
We addressed the portfolio optimization problem using the SCIP optimizer, treating its solution as the benchmark reference (see Supplementary Material). The running time was determined by averaging the results from $100$ executions. This average running time served as the basis for comparing the performance of classical methods with hybrid VQE routines. However, it is important to note that we did not expect that the VQE approach would outperform the classical methods in terms of speed.
The experiments were performed in three different scenarios as described in Table 2. For Scenario I, we ran the VQE algorithm for $20$ different ansatzes (Dicke state, $R_{y}$, and $18$ variations of Two Local). For Scenarios II and III only the Dicke state was used. For all scenarios, we tested $5$ classical optimizers with $1000$ iterations, $100$ randomly sampled initial points (ansatz parameters), $4096$ shots per circuit, totalizing more than $60$ billion executions ($\approx n_{ansatz}\times n_{optimizers}\times n_{executions}\times n_{shots}\times n_{iterations}$) . More details of the variations of the two-local ansatz used here can be found in Supplementary Material. All data used in this work are publicly available and were obtained using the Yahoo Finance API.
@@
From a theoretical standpoint, the improved optimization performance of the Dicke-state ansatz arises from its structural design, which confines the variational search to the feasible subspace defined by the diversification constraints. By preparing the quantum state as a tensor product of Dicke states, the search space is reduced from the full $2^{n}$ Hilbert space to $\prod_{i}C_{n_{i},k_{i}}$ a substantial reduction in realistic scenarios where $k_{i}\ll n_{i}$. This restriction removes infeasible configurations, concentrates the optimization on meaningful portfolio states, and preserves the symmetry of fixed Hamming weight. Consequently, the optimizer operates on a smoother energy landscape, improving convergence stability and accuracy relative to more general ansatzes.
Scenario $n_{a}$ $n_{c}$ $n_{s}$ Ansatz $n_{p}$ $n_{\text{search space}}$
I $10$ 11 44 $|D_{4}^{10}\rangle$ $30$ $210$
II $25$ 55 55 $|D_{1}^{5}\rangle^{\otimes\ 5}$ 2020 $3125$
III $25$ 55 99 $|D_{2}^{5}\rangle^{\otimes\ 2}|D_{1}^{5}\rangle^{\otimes 2}|D_{3}^{5}\rangle$ $31$ $25000$
Table 2: The table shows the different scenarios we use to evaluate the performance of parametrized Dicke state in the context of multiclass portfolio optimization. Where $n_{a}$ represents the number of assets (that define the number of qubits used in the scenario), $n_{c}$ number of classes, $n_{s}$ amount of select assets, $n_{p}$ number of parameters of Dicke state ansatz. $|D_{k}^{n}\rangle$ is a superposition of all states with $n$ qubits and $k$ qubits equal $|1\rangle$. Each state represents a class where $n$ is the total number of assets available to choose and $k$ is the number of assets we must select from it. The parameter $\vec{\theta}$ was omitted in the state for better visualization. The last column shows the actual size of search space for each scenario.
@@
Figure 5: Plots $(a)$, $(b)$, and $(c)$ compare the approximation ratio ($a_{r}$) distributions before (gray) and after (colored) optimization for each scenario (columns) and optimizers (rows). The gray represent $a_{r}$ distributions with the initial parameters, while the colored ones reflect optimized parameters, based on $100$ experiments per optimizer. A vertical black dashed line marks $a_{r}=0.9$. Across all optimizers and scenarios, the initial distributions shift toward $a_{r}\geq 0.9$, indicating the positive impact of the hybrid approach and optimizers.In Scenario I (blue) all the optimizers guide the solution to the optimal region.
@@
Another challenge in the evaluation of hybrid algorithms is to measure the effect of the classical optimizer on the solution. As mentioned before, we tested five different optimizers: CMA-ES [grayver2016exploring], COBYLA [powell1994direct], Random Sampler, SPSA [spall1998overview] and QNSPSA [gacon2021simultaneous]. The evaluation takes into account three different metrics: the approximation ratio, the frequency with which the right answer appears in the $100$ trials per optimizer, and the quality of the quantum output for each trial (the probability of measuring the target state). The approximation ratio is defined as [shaydulin2019evaluating, ZhouLeo2020Quantum, herrman2022multi, abbas2024challenges]
@@
where $\langle H\rangle_{\psi(\vec{\theta^{*})}}$ is the expected value of the problem Hamiltonian computed with the ansatz and the set of parameters obtained at the end of optimization. $E_{0}$ is the ground-state energy associated with the optimal solution that was calculated by computing the lowest eigenvalue for the problem Hamiltonian. $E_{max}$ represents the highest Hamiltonian eigenvalue. The metric above measures how close or distant the solution is from the optimal. For example, an approximation ratio equal to $1$ means that the solution is equal to the optimal one.